Define a partition of an interval. Write any three different examples of

partitions of [0,1].

(ii). Let a real valued function 𝑓 be defined and bounded over [𝑎, 𝑏]. Prove that 𝑓 is

Riemann integrable over [𝑎, 𝑏] if for each 𝜖 > 0 there is a partition 𝑃 such that

𝑈(𝑃, 𝑓) − 𝐿(𝑃, 𝑓) < 𝜖

(iii). Show that 𝑓(𝑥) = 𝑥2

is Riemann integrable over [0,2].


1
Expert's answer
2021-05-31T15:12:00-0400

Ans:-

"(i)" Partition of a Set is defined as "A collection of disjoint subsets of a given set. The union of the subsets must equal the entire original set".

three different examples of partitions of "[0,1]" are "(0,0.4) \ , \ (0.4,0.6)\ , \ (0.6,1)"


"(ii)" Let f  be bounded on "[a,b]" . Then f  is Riemann integrable if and only if for every  ϵ there is a partition on "[a,b]"  such that: "0\u2264U(f,P)\u2212L(f,P)\u2264\u03f5"


Since

"\int_ a ^b f=inf[{U(f,P)}]=sup[{L(f,P)}]"


there should exist two partitions "P_1 \ and \ P_2"  such that:


"0≤U(f,P_1)−\int _a ^b f<\dfrac{ϵ}2"


and

"0≤\int^b_af−L(f,P_2)<\dfracϵ2"


The argument is almost identical for "L(f,P)" (except flipping some inequalities)

Hence the function "f" is Riemann integrable


"(iii) \ f(x)= x^2"

For Reimann integrable


"\Rightarrow ∫^b _a f(x)dx"

So We have this expression equal to

"\Rightarrow ∫^b _0 x^2dx−∫^a _0 x^2dx"


"\Rightarrow (b^3−a^3)∫^1_0 x^2dx"


Here "a=0 \ \ and \ \ b=2"

Hence "\ud835\udc53(\ud835\udc65) = \ud835\udc65^2" is Riemann integrable over "[0,2]"